Find the real solutions of teh system of equations:
x^2 + y^2 = 4
2x + y = 0
Thanks!

Answer by richwmiller(17219)

y=-2x
x^2+(-2x)^2=4
x^2+4x^2=4
5x^2=4
x^2=4/5
x=+-2/sqrt(5)
x=+2sqrt(5)/5
x=-2sqrt(5)/5
y=-2x